A car starts from rest and accelerates uniformly over a time of 5.21 seconds for a distance of 110 m. Determine the acceleration of the car.


1
Expert's answer
2021-07-02T09:38:38-0400

Given:

"v_i=0\:\rm m/s"

"t=5.21\:\rm s"

"d=110\:\rm m"

The distance

"d=v_it+\frac{at^2}{2}"

Hence,

"a=\frac{2d}{t^2}=\frac{2*110\:\rm m}{(5.21\;\rm s)^2}=8.10\:\rm m/s^2"